# Bright and dark breathers on an elliptic wave in the defocusing mKdV equation

### v1

@@ 5.2. Rational solutions
It follows from (1.7) with expressions for $b,c,d$ that $\zeta_{1}=-v^{-1}$ and $\zeta_{2}=\zeta_{3}=v^{-1}$. By using parameter $a\in\mathbb{C}$, we get from (3.6) and (3.16) that
@@
from which we obtain two pairs of roots labeled as $\pm y_{+}$ and $\pm y_{-}$, where
$y_{\pm}=\sqrt{1+\frac{4}{\pm\sqrt{1+2(1-z^{2})(1+iz^{-1}\sqrt{1-z^{2}})}-1}}.$ (5.31)
Similarly, we obtain from (5.2) an equation
@@
with two pairs of roots labeled as $\pm y_{+}^{*}$ and $\pm y_{-}^{*}$, where
$y_{\pm}^{*}=\sqrt{1+\frac{4}{\pm\sqrt{1+2(1-z^{2})(1-iz^{-1}\sqrt{1-z^{2}})}-1}}.$ (5.33)
Expressions (5.31) and (5.33) have square root singularities at $z=0$ and $z=\pm 1$.
@@
obtained from (5.4) with the scaling (5.29). Hence, based on the numerical data and the scaled transformation (5.29), we have
$x_{1}=-\frac{v}{2}y_{+},\quad x_{2}=-\frac{v}{2}y_{-},\quad x_{1}^{*}=\frac{v}{2}y_{+}^{*},\quad x_{2}^{*}=\frac{v}{2}y_{-}^{*}.$
We also note that $y_{\pm}^{*}=\overline{y}_{\pm}$, which suggests that $x_{1,2}^{*}=-\overline{x}_{1,2}$.
@@ 5.3. Hyperbolic solutions
If $0<\zeta_{3}=\zeta_{2}<\zeta_{1}$, we have $k=1$ from (1.9). Since $K=\infty$ and $K^{\prime}=\frac{\pi}{2}$, the fundamental rectangle $K,K]\times[-iK^{\prime},iK^{\prime}]$ becomes a horizontal strip of the width $\pi$ in the vertical direction. Equation (5.6) becomes
$4i\zeta(\zeta_{1}^{2}-\zeta_{2}^{2})\frac{\zeta_{2}^{2}\cosh^{2}(2z)+\zeta_{1}^{2}-2\zeta_{2}^{2}}{(\zeta_{1}+\zeta_{2}\cosh(2z))^{2}}+4i\zeta(\zeta^{2}-\zeta_{1}^{2})+\mu=0,$
which can be expanded as
@@
Since $\mu^{2}=-16(\zeta^{2}-\zeta_{1}^{2})(\zeta^{2}-\zeta_{2}^{2})^{2}$, multiplying by $\left(\zeta^{2}-\zeta_{2}^{2}-\frac{\mu}{4i\zeta}\right)$ yields after straightforward computations:
@@
We consider $\zeta\in(0,\zeta_{2})$ and select $\mu=4(\zeta_{2}^{2}-\zeta^{2})\sqrt{\zeta_{1}^{2}-\zeta^{2}}$ from (3.19). Solving the previous equation for $\cosh(2z)$ yields two solutions:
$\cosh(2z)=\frac{-\zeta_{2}^{2}(\zeta_{1}^{2}-\zeta^{2})+i\zeta(\zeta_{1}^{2}-\zeta_{2}^{2})\sqrt{\zeta_{1}^{2}-\zeta^{2}}\pm(\zeta_{1}^{2}-\zeta_{2}^{2})\sqrt{\zeta^{2}(2\zeta_{2}^{2}-\zeta_{1}^{2})-2i\zeta\zeta_{2}^{2}\sqrt{\zeta_{1}^{2}-\zeta^{2}}}}{\zeta_{1}\zeta_{2}(\zeta_{2}^{2}-\zeta^{2})}.$ (5.36)
Roots of (5.36) solve equation (5.1). Hence, they are denoted by $\pm z_{1}$ and $\pm z_{2}$ in $(-\infty,\infty)\times\left[-\frac{i\pi}{2},\frac{i\pi}{2}\right]$. Similarly, roots of
solve equation (5.2). Hence, they are denoted by $\pm z_{1}^{*}$ and $\pm z_{2}^{*}$ in $(-\infty,\infty)\times\left[-\frac{i\pi}{2},\frac{i\pi}{2}\right]$. We detect numerically which of the four roots of (5.36) and (5.37) are simultaneously roots of (5.3). They are labeled as $z_{1}$, $z_{2}$, $z_{1}^{*}$, $z_{2}^{*}$. The sign-reflected roots $z_{1}$, $z_{2}$, $z_{1}^{*}$, $z_{2}^{*}$ of (5.36) and (5.37) are simultaneously roots of (5.4). The correspondence between the roots of $z$ and roots of $x$ is $z=\nu x$, where $\nu=\sqrt{\zeta_{1}^{2}-\zeta_{2}^{2}}$.
Figure 10 displays roots $z_{1}$, $z_{2}$, $z_{1}^{*}$, $z_{2}^{*}$ of (5.36) and (5.37) for $\zeta_{1}=1$ and $\zeta_{2}=\zeta_{3}=0.5$. The roots are parameterized by $\zeta\in(0,\zeta_{2})$. The big dots of the same color show roots at $\zeta=\zeta_{2}$ and $\zeta=0$. Dashed horizontal lines in magenta shows the boundary of the vertical strip at $\pm iK^{\prime}=\pm\frac{\pi i}{2}$. For $\zeta=0$, four pairs of roots are located at $\pm iK^{\prime}\pm\alpha$. As $\zeta\to\zeta_{2}$, four roots diverge to infinity since $K=\infty$ and two pairs of roots coalesce on $i\mathbb{R}$.
The solid lines show the roots of (5.1) and (5.2) which are simultaneously roots of (5.3), that is, roots labeled as $\{z_{1},z_{2},z_{1}^{*},z_{2}^{*}\}$. Consequently, the dashed lines show the reflected roots $\{-z_{1},-z_{2},-z_{1}^{*},-z_{2}^{*}\}$ which are roots of (5.1) and (5.2) which are simultaneously roots of (5.4). We can see that the roots satisfy the symmetry $z_{1,2}^{*}=-\bar{z}_{1,2}$, as in the case of rational solutions.
@@
Acknowledgement. The first author thanks M. Bertola, D. Korotkin, and L. Ling for many discussions related to this project. The main breakthrough for this project was obtained during the Third Joint Alabama-Florida Conference on Differential Equations, Dynamical Systems and Applications at University of Alabama Birmingham (UAB), May 20-22, 2025. Some computations in Section 5 were obtained earlier, in collaboration with L. K. Arruda during the previous project [3].

### v2

@@ 5.2. Rational solutions
It follows from (1.7) with expressions for $b,c,d$ that $\zeta_{1}=-v^{-1}$ and $\zeta_{2}=\zeta_{3}=v^{-1}$. By using parameter $a\in\mathbb{C}$, we get from $\zeta^{2}=\wp(v)-\wp(a)$ and $\mu=-2\wp^{\prime}(a)$ that
@@
The following lemma defines the roots $\{x_{1},x_{2}\}$ and $\{x_{1}^{*},x_{2}^{*}\}$ of Lemma 7 and 8 for the rational solutions.
Lemma 10.
The roots $\{x_{1},x_{2}\}$ and $\{x_{1}^{*},x_{2}^{*}\}$ of (5.1) and (5.2) for the rational solution (5.28) are given by $x_{1,2}=\frac{v}{2}y_{\pm}$ and $x_{1,2}^{*}=\frac{v}{2}y_{\pm}^{*}$, where $\{y_{+},y_{-}\}$ and $\{y_{+}^{*},y_{-}^{*}\}$ are defined up to the sign choice by
$y_{\pm}=\sqrt{1+\frac{4}{\pm\sqrt{1+2(1-z^{2})(1+iz^{-1}\sqrt{1-z^{2}})}-1}}$ (5.29)
and
$y_{\pm}^{*}=\sqrt{1+\frac{4}{\pm\sqrt{1+2(1-z^{2})(1-iz^{-1}\sqrt{1-z^{2}})}-1}},$ (5.30)
where $z=v\zeta$.
Proof.
@@
from which we obtain two pairs of roots $\{\pm y_{+},\pm y_{-}\}$ in the form (5.29). Similarly, equation (5.2) can be rewritten in the form
@@
from which we obtain two pairs of roots $\{\pm y_{+}^{*},\pm y_{-}^{*}\}$ in the form (5.30). ∎
@@
obtained from (5.4) with the scaling (5.32). The appropriate roots of (5.29) and (5.30), which are simultaneously roots of (5.35) and (5.36) are identified with the help of computer computations.
Thus, based on the numerical data and the scaling transformation (5.32), we have
$x_{1}=-\frac{v}{2}y_{+},\quad x_{2}=-\frac{v}{2}y_{-},\quad x_{1}^{*}=\frac{v}{2}y_{+}^{*},\quad x_{2}^{*}=\frac{v}{2}y_{-}^{*},$
from which the relation $y_{\pm}^{*}=\overline{y}_{\pm}$ suggests that $x_{1,2}^{*}=-\overline{x}_{1,2}$. We emphasize that the roots $\{x_{+},x_{-}\}$ and $\{x_{+}^{*},x_{-}^{*}\}$ obtained from (5.29) and (5.30) depends on the spectral parameter $\zeta$ via the square root singularities at $z=0$ and $z=\pm 1$.
@@ 5.3. Hyperbolic solutions
If $0<\zeta_{3}=\zeta_{2}<\zeta_{1}$, we have $k=1$ from (1.9), which corresponds to the solitary wave rather than the periodic wave of the mKdV equation (1.1). Since $K=\infty$ and $K^{\prime}=\frac{\pi}{2}$, the fundamental rectangle $K,K]\times[-iK^{\prime},iK^{\prime}]$ becomes a horizontal strip of the width $\pi$ in the vertical direction.
The following lemma defines the roots $\{x_{1},x_{2}\}$ and $\{x_{1}^{*},x_{2}^{*}\}$ of Lemma 7 and 8 for the hyperbolic solutions.
Lemma 11.
The roots $\{x_{1},x_{2}\}$ and $\{x_{1}^{*},x_{2}^{*}\}$ of (5.1) and (5.2) for the hyperbolic solution (1.8) with $0<\zeta_{3}=\zeta_{2}<\zeta_{1}$ are given by $x_{1,2}=\nu^{-1}z_{1,2}$ and $x_{1,2}^{*}=\nu^{-1}z_{1,2}^{*}$, where $\{z_{1},z_{2}\}\in\mathbb{R}\times\left[-\frac{i\pi}{2},\frac{i\pi}{2}\right]$ and $\{z_{1}^{*},z_{2}^{*}\}\in\mathbb{R}\times\left[-\frac{i\pi}{2},\frac{i\pi}{2}\right]$ are defined up to the sign choice by roots of
$\cosh(2z)=\frac{-\zeta_{2}^{2}(\zeta_{1}^{2}-\zeta^{2})+i\zeta(\zeta_{1}^{2}-\zeta_{2}^{2})\sqrt{\zeta_{1}^{2}-\zeta^{2}}\pm(\zeta_{1}^{2}-\zeta_{2}^{2})\sqrt{\zeta^{2}(2\zeta_{2}^{2}-\zeta_{1}^{2})-2i\zeta\zeta_{2}^{2}\sqrt{\zeta_{1}^{2}-\zeta^{2}}}}{\zeta_{1}\zeta_{2}(\zeta_{2}^{2}-\zeta^{2})}$ (5.37)
and
where $\zeta\in(0,\zeta_{2})$.
Proof.
It follows from the proof of Lemma 7 that equation (5.1) can be written as equation (5.6), which is further rewritten as
$4i\zeta(\zeta_{1}^{2}-\zeta_{2}^{2})\frac{\zeta_{2}^{2}\cosh^{2}(2z)+\zeta_{1}^{2}-2\zeta_{2}^{2}}{(\zeta_{1}+\zeta_{2}\cosh(2z))^{2}}+4i\zeta(\zeta^{2}-\zeta_{1}^{2})+\mu=0.$
This equation can be expanded as
@@
Since $\mu^{2}=-16(\zeta^{2}-\zeta_{1}^{2})(\zeta^{2}-\zeta_{2}^{2})^{2}$, multiplying by $\left(\zeta^{2}-\zeta_{2}^{2}-\frac{\mu}{4i\zeta}\right)$ yields the following equation after straightforward computations:
@@
If $\zeta\in(0,\zeta_{2})$, then $\mu=4(\zeta_{2}^{2}-\zeta^{2})\sqrt{\zeta_{1}^{2}-\zeta^{2}}$ follows from (3.19). Solving the previous equation for $\cosh(2z)$ yields two solutions given by (5.37). Similarly, equation (5.2) can be solved with two solutions given by (5.38). ∎
@@
Figure 10 displays roots $\{\pm z_{1},\pm z_{2}\}$ and $\{\pm z_{1}^{*},\pm z_{2}^{*}\}$ of (5.37) and (5.38) for $\zeta_{1}=1$ and $\zeta_{2}=\zeta_{3}=0.5$. The roots are parameterized by $\zeta\in(0,\zeta_{2})$. The big dots of the same color show roots at $\zeta=\zeta_{2}$ and $\zeta=0$. Dashed horizontal lines in magenta shows the boundary of the vertical strip at $\pm iK^{\prime}=\pm\frac{\pi i}{2}$. For $\zeta=0$, four pairs of roots are located at $\pm iK^{\prime}\pm\alpha$. As $\zeta\to\zeta_{2}$, four roots diverge to infinity since $K=\infty$ and two pairs of roots coalesce on $i\mathbb{R}$.
The solid lines show the roots of (5.37) and (5.38) which are simultaneously roots of (5.3), that is, roots labeled as $\{z_{1},z_{2},z_{1}^{*},z_{2}^{*}\}$. Consequently, the dashed lines show the reflected roots $\{-z_{1},-z_{2},-z_{1}^{*},-z_{2}^{*}\}$ which are roots of (5.37) and (5.38) which are simultaneously roots of (5.4). Again, the roots are identified with the help of computer computations. We can see from Figure 10 that the roots satisfy the symmetry $z_{1,2}^{*}=-\bar{z}_{1,2}$, which suggests the same relation $x_{1,2}^{*}=-\overline{x}_{1,2}$ as in the case of rational solutions.
Acknowledgement. The first author thanks M. Bertola, D. Korotkin, and L. Ling for many discussions related to this project. The main breakthrough for this project was obtained during the Third Joint Alabama-Florida Conference on Differential Equations, Dynamical Systems and Applications at University of Alabama Birmingham (UAB), May 20-22, 2025. Some computations in Section 5 were obtained earlier, in collaboration with L. K. Arruda during the previous project [3]. The work of D. E. Pelinovsky was supported in part by the National Natural Science Foundation of China (grant no. 12371248).